The [a] of pure quinine, an antimalarial drug, is -165. If a solution contains 72% quinine and 28% of its enantiomer (ee=44%), what is [a] for the solution?
Answer #1

a of solution=0.44*(-165)

=-72.6
Answer #2
net rotation = 0.72-0.28=0.44 so [a] for solution = -0.44*165 = -72.6 ans ans -72.6